You run 100 m in 12 s, then turn around and jog 50 m back toward the starting point in 30 s.
a. Calculate your average speed.
1
Expert's answer
2019-10-22T10:18:51-0400

Average speed is given by:


"v_{aver}=\frac{S_{total}}{t_{total}}=\frac{S_1+S_2}{t_1+t_2}=\frac{100+50}{12+30}=4.7 m/s"